OpenStudy (anonymous):

Rewrite with only sin x and cos x. sin (3x) - cos (x) Please guide me with showing work because the process of substituting is what I'm having difficulties with.

OpenStudy (anonymous):

So do you have any idea how to start?

OpenStudy (anonymous):

sin (2x + x) - cos x ?

OpenStudy (anonymous):

The identity for Sin(3x) is sin(3x) = -sin^3(x) + 3cos^2(x)sin(x)

OpenStudy (anonymous):

I'm lost? These are the choices:

OpenStudy (anonymous):

substitute the identity I gave you into the problem. This gives -sin3(x)+3cos2(x)sin(x)-cos(x). Rearrange for positive term convention. 3cos2(x)sin(x-sin3(x)-cos(x) is the answer

OpenStudy (anonymous):

Don't see it in the answer choices...

OpenStudy (anonymous):

you're right, but the other identity for sin(3x) is -4sin3(x)+3sin(x). That definitly isn't the route to an answer.

OpenStudy (campbell_st):

you need to start by rewriting sin(3x) sin(3x) = sin(2x + x) = sin(2x)cos(x) + cos(2x)sin(x) you now need sin(2x) = 2sin(x)cos(x) and cos(2x) = sos^2(x) - sin^2(x) and substitute them so you will have sin(3x) - cos(x) = sin(2x)cos(x) + cos(2x)sin(x) - cos(x) = 2sin(x)cos(x)cos(x) + (cos^2(x) - sin^2(x))sin(x) - cos(x) = 2 sin(x)cos^2(x) + sin(x)cos^2(x) - sin^3(x) - cos(x) I'll let you finish by collecting like terms...

OpenStudy (anonymous):

Omg, I understood it... Thank you so much! :)